Using the Pythagorean Theorem, find the length of a leg of a right triangle if the other leg is 8 feet long and the hypotenuse is 10 feet long.
Answer by cleomenius(959)

8^2 + x^2 = 10^2
64 + x^2 = 100
x^2 = 36
x = 6 feet
Cleomenius.